# StableToolBench: Towards Stable Large-Scale Benchmarking on Tool Learning of Large Language Models

### v2

@@ Abstract
[…]
@@ 1 Introduction
With the rapid developments of Large Language Models (LLMs; Brown et al., 2020; Gemini Team, 2023; OpenAI, 2023; Touvron et al., 2023), tool learning which leverage LLMs to schedule a variety of external tools has attracted enormous attention Nakano et al. (2022); Yao et al. (2023); Lu et al. (2023). Project page: https://zhichengg.github.io/stb.github.io/GitHub page: zhichengg/StableToolBench Previous studies (Hao et al., 2023; Hsieh et al., 2023; Schick et al., 2023; Tang et al., 2023) aim to augment LLMs with tools to enhance performance on conventional natural language processing (NLP) downstream tasks, while recent work (Qin et al., 2023a; Yao et al., 2022; Cai et al., 2024) primarily focus on solving real-world scenarios that require the use of tools. In general, tool learning complements the capabilities of vanilla LLMs and bridges the gap to real-world applications.
@@
To assess the capability of LLMs to use tools, a series of tool learning benchmarks have been introduced. Several pioneering studies have heavily relied on human-crafted offline tools Yang et al. (2023b); Xu et al. (2023) or hand-selected online tools Li et al. (2023a, b); Chen et al. (2023b). While these tools are high-quality, their scale remains relatively small, thereby limiting their ability to accurately reflect real-world scenarios. To address this limitation, subsequent studies (Tang et al., 2023; Ye et al., 2024; Qin et al., 2023b) have advocated for leveraging extensive collections of online tools that span across various domains. Owing to the increased scale, the automatic evaluation of tool learning has moved closer to real-world scenarios.
@@
However, concerns have been raised regarding the reproducibility and comparability of benchmark performance over time. Firstly, due to the difficulty of complex tasks with tool usage, the common automatic evaluator gpt-3.5 can not assume the role of discriminating. The automatic evaluator cannot determine whether the task is solvable or unsolvable as discussed in Section 2.2, thus the capability limitation introduces randomness to model performances. Secondly, the stability of API status in more than half of online tools (55.6% in ToolBench) is fluctuant. On the one hand, users may be requested to authorise to use of these tools or APIs. On the other hand, these online tools provided by developers may be available during the initial construction of the benchmark but become inaccessible later. This leads to a problem where the constructed queries in the benchmarks may no longer be completed with their original referenced tools. Therefore, it is crucial to strike a balance between enhancing the stability of these benchmarks and maintaining their diversity and scope.
@@
To this end, we propose a new benchmark, named StableToolBench, which proposes a virtual API system and a stable evaluation system. In the virtual API system, we first build a caching system to store the output of API calls. This ensures the stability and reproducibility of API behaviours. Given that the benchmark questions are limited, our caching system can cover a significant number of API call scenarios. However, merely using a cache is not enough because there still exists a large number of unavailable APIs. To resolve this problem, we use LLMs to simulate the behaviours of these APIs. Specifically, we feed the documentation and few-shot real API calls if available in the cache to LLMs and ask LLMs to mock the behaviour of the APIs given a request. As a result, users can always get responses from APIs in an indistinguishable way as long as the LLMs are accessible. On the whole, our system first tries to find a hit in the cache. Unless there is a cache miss and a real API call is not received, the simulated server will be used. As for the stable evaluation system, we design two metrics (i.e., SoPR and SoWR) after judging solvable tasks and replace all the automatic evaluators with gpt-4 to mitigate the randomness and indistinguishability during evaluation. Experiments demonstrate that our virtual API system consistently maintains stability in the face of real API modifications. Furthermore, our evaluation system yields more stable results and exhibits a high level of concordance with human assessments.
@@
The main contributions of our work are summarized as follows:
@@
  • Extensive experiments show that our benchmark provides much more stable model performance, robust to various types API failures.
  • In addition to enhancing stability, our virtual API system exhibits realism and diversity comparable to that of the real API system.
@@ 3.1 Virtual API Server
3.5 0613 (C) 55.9±$\pm{1.0}$ 50.8±$\pm{0.8}$ 55.9±$\pm{1.0}$ 44.1±$\pm{0.8}$ 36.2±$\pm{0.4}$ $\pm{1.5}$ 49.1±$\pm{1.0}$
3.5 0613 (D) 66.4±$\pm{1.5}$ 64.3±$\pm{1.0}$ 67.2±$\pm{2.4}$ 67.7±$\pm{0.8}$ 61.5±$\pm{1.0}$ 81.4±$\pm{1.5}$ 68.1±$\pm{1.4}$
4 0613 (C) 50.7±$\pm{0.4}$ 57.1±$\pm{0.3}$ 51.9±$\pm{0.3}$ 55.0±$\pm{1.1}$ 61.6±$\pm{0.8}$ 56.3±$\pm{0.8}$ 55.4±$\pm{0.6}$
4 0613 (D) 65.5±$\pm{1.1}$ 62.0±$\pm{1.7}$ 72.1±$\pm{1.6}$ 70.8±$\pm{1.3}$ 73.1±$\pm{1.4}$ 74.9±$\pm{1.5}$ 69.7±$\pm{1.4}$
T-LLaMA (C) 37.2±$\pm{0.1}$ 42.3±$\pm{0.4}$ 43.0±$\pm{0.5}$ 37.4±$\pm{0.4}$ 33.6±$\pm{1.2}$ 39.6±$\pm{1.0}$ 38.9±$\pm{0.6}$
T-LLaMA (D) 59.8±$\pm{1.5}$ 59.5±$\pm{1.4}$ 65.7±$\pm{1.1}$ 56.5±$\pm{0.3}$ 47.6±$\pm{0.4}$ 62.8±$\pm{1.9}$ 58.7±$\pm{1.1}$
3.5 1106 (C) 51.3±$\pm{0.6}$ 48.8±$\pm{0.3}$ 59.9±$\pm{0.8}$ 50.8±$\pm{0.7}$ 43.2±$\pm{0.8}$ 58.5±$\pm{0.8}$ 52.1±$\pm{0.7}$
3.5 1106 (D) 67.8±$\pm{0.9}$ 67.2±$\pm{0.3}$ 72.9±$\pm{0.7}$ 63.2±$\pm{1.0}$ 70.9±$\pm{0.4}$ 77.6±$\pm{0.8}$ 69.9±$\pm{0.7}$
4 Turbo (C) 63.1±$\pm{1.0}$ 64.5±$\pm{0.5}$ 55.3±$\pm{0.3}$ 63.0±$\pm{0.8}$ 57.3±$\pm{0.8}$ 61.7±$\pm{0.8}$ 60.8±$\pm{0.7}$
4 Turbo (D) 70.8±$\pm{1.0}$ 71.1±$\pm{0.7}$ 70.4±$\pm{1.2}$ 70.4±$\pm{1.3}$ 71.7±$\pm{0.4}$ 84.7±$\pm{1.7}$ 73.2±$\pm{1.1}$
Table 3: Solvable pass rate scores. We run all models once, evaluate three times and take the average results. “3.5 0613”, “4 0613”, “3.5 1106”, “4 Trubo”, “T-LLaMA” stands for gpt-3.5-turbo-0613, gpt-4-0613, gpt-3.5-turbo-1106, gpt-4-turbo-preview, ToolLLaMA v2 respectively. C and D stand for CoT and DFS respectively. The experiments below follow the denotation.
@@ 3.2 Stable Evaluation System
We then report Solvable Pass Rate (SoPR) and Solvable Win Rate (SoWR) based on our obtained solvable tasks. Due to the limitation of gpt-3.5-turbo-16k in tool learning, we uniformly adopt gpt-4-turbo-preview as the automatic evaluator. SoPR is in essence PR with all tasks solvable and only assesses the answers using the same prompt in ToolBench. The evaluator assigns outcomes of answers categorised as Solved, Unsolved, or Unsure, which respectively contribute scores of 1, 0.5, and 0 to the overall SoPR calculation. As for SoWR, when one is solved and the other is unsolved, the solved one wins. Under other circumstances, gpt-4-turbo-preview will be used to make a win-lose decision.
@@ 4.1 Performance
3.5 0613 (D) 57.7 60.8 61.4 63.2 66.1 70.5 63.3
4 0613 (C) 50.3 54.2 50.6 64.2 50.0 55.7 54.2
4 0613 (D) 57.1 60.1 57.0 74.5 64.5 72.1 64.2
T-LLaMA (C) 35.0 30.7 37.3 36.8 31.5 23.0 32.4
T-LLaMA (D) 43.6 45.1 38.6 53.8 42.7 45.9 44.9
3.5 1106 (C) 46.6 45.1 48.1 37.7 44.4 52.5 45.7
3.5 1106 (D) 56.4 54.2 51.9 62.3 54.0 72.1 58.5
4 Turbo (C) 68.7 71.9 58.2 76.4 71.0 73.8 70.0
4 Turbo (D) 66.9 73.9 68.4 78.3 72.6 77.0 72.9
Table 5: Solvable Win Rate scores. We run all models once against GPT-3.5-Turbo-0613 + CoT and evaluate three times. We follow the ToolBench implementation to take the most frequent result for each query during evaluation.
@@ 4.2 Stability of Virtual API Server
Figure 7: Performance change when manually making APIs down with our virtual online API system. The results are averaged over all six groups. Solving rates are reported. We run each experiment one time and evaluate it three times and take the average score.
@@ 4.5 Effectiveness of Caching System
Table 7: Cache hit rate (%) with various models and methods. Final, Mid, Start represent the final version of the cache, the mid-way version containing 151,152 (91.6%) items of the final version, and the starting version containing only train and test set. Experiments are independent runs of Section 4.1 with fixed cache, run on 13 Feb 2024.
4.6 Human Evaluation of Evaluator
Considering that GPT-3.5 is limited to evaluating the performance in tool learning, we replace the automatic evaluators with stronger LLMs. In this section, we manually assess the correctness of different automatic evaluators. Here, we sample 100 task solvable questions, 50 answer-solving questions in the PS / SoPR evaluation, and 50 comparison questions in the WR / SoWR evaluation from the experiments running. We then collect all the corresponding answers of different LLMs during previous evaluations. These questions are further manually labeled by three human annotators to obtain the ground truth. With the ground truth, we calculate the accuracy scores of these models and show the results in Table 8. It can be seen that our used LLMs (i.e., Claude 2, Gemini, and GPT-4) are much better than GPT-3.5 in ToolBench to determine the solvability of tasks, where the Gemini and GPT-4 outperform by a large margin. In both the evaluation of answers and comparison, GPT-4 significantly outperforms GPT-3.5, especially in the comparison to compute WR. It is worth noting that all the accuracies of GPT-3.5 are lower than 70%, indicating that GPT-3.5 cannot assess the performance in tool learning.
@@ 5 Related Work
Recent studies have shed light on the burgeoning capabilities of LLMs in understanding and mastering tools (Li et al., 2023b; Patil et al., 2023; Yang et al., 2023b; Song et al., 2023; Tang et al., 2023; Ye et al., 2024; Xu et al., 2023). Gaining access to external tools endows LLMs with real-time factual knowledge (Yang et al., 2023a), multimodal functionalities (Gupta and Kembhavi, 2023), and specialized skills in vertical domains (Jin et al., 2023). However, few work has been done to explore the stability of the tool environment in specific benchmarks and how it affects the LLMs performance in tool-augmented tasks.
@@
Evaluating the performance of LLMs in tool-augmented tasks presents unique challenges and opportunities. Numerous works have been developed for the assessment of tool utilisation, primarily emphasising response comparison  (Zhuang et al., 2023), tool call accuracy  (Patil et al., 2023), or a synthesis of these aspects  (Li et al., 2023b). Distinguishing itself, Qin et al. (2023c) introduces an innovative methodology by integrating a large language model (LLM) as a judge to evaluate the comprehensive solution path. Subsequent research  (Wang et al., 2023) focus on the multi-turn interaction capabilities of LLMs with both tools and user feedback. In a departure from the aforementioned approaches, Chen et al. (2023b) presents itself as the inaugural benchmark specifically tailored for the fine-grained assessment of tool utilisation capabilities. However, there exists a gap in the literature concerning the exploration of evaluation stability when evaluating the tool usage capabilities of LLMs.

### v4

@@ Abstract
Zhicheng Guo1,2121,21 , 2, Sijie Cheng1,2,31231,2,31 , 2 , 3, Hao Wang4444, Shihao Liang5555, Yujia Qin1111, Peng Li2222, Zhiyuan Liu1111, Maosong Sun1111, Yang Liu1,2,6 $1$Dept. of Comp. Sci. & Tech., Institute for AI, Tsinghua University, Beijing, China $2$Institute for AI Industry Research (AIR), Tsinghua University, Beijing, China 333301.AI $4$Google $5$The University of Hong Kong $6$Jiangsu Collaborative Innovation Center for Language Competence, Jiangsu, China {guo-zc21, csj23}@mails.tsinghua.edu.cn
@@ 1 Introduction
With the rapid developments of Large Language Models (LLMs; Brown et al., 2020; Gemini Team, 2023; OpenAI, 2023; Touvron et al., 2023), tool learning which leverage LLMs to schedule a variety of external tools has attracted enormous attention Nakano et al. (2022); Yao et al. (2022b); Lu et al. (2023). Project: zhichengg.github.io/stb.github.io/GitHub: THUNLP-MT/StableToolBench Previous studies (Hao et al., 2023; Hsieh et al., 2023; Schick et al., 2023; Tang et al., 2023) aim to augment LLMs with tools to enhance performance on conventional natural language processing (NLP) downstream tasks, while recent work (Qin et al., 2023a; Yao et al., 2022a; Cai et al., 2024) primarily focus on solving real-world scenarios that require the use of tools. In general, tool learning complements the capabilities of vanilla LLMs and bridges the gap to real-world applications.
@@
To assess the capability of LLMs to use tools, a series of tool learning benchmarks have been introduced. Several pioneering studies have heavily relied on human-crafted offline tools Yang et al. (2023b); Xu et al. (2023) or hand-selected online tools Li et al. (2023b, a); Chen et al. (2023b). While these tools are high-quality, their scale remains relatively small, thereby limiting their ability to accurately reflect real-world scenarios. To address this limitation, subsequent studies (Tang et al., 2023; Ye et al., 2024; Qin et al., 2023b) have advocated for leveraging extensive collections of online tools that span across various domains. Owing to the increased scale, the automatic evaluation of tool learning has moved closer to real-world scenarios. However, concerns have been raised regarding the stability of these online tools, which has implications for the reproducibility and comparability of benchmark performance over time [^1] According to its definition, benchmarks should remain stable, and the model performance assessed on them must be comparable over time.. For instance, the well-recognised ToolBench [^2] We use ToolEval2 in ToolBench as the benchmark. Qin et al. (2023c) has shown performance discrepancies that cannot be reproduced months after its release, as analysed in  Section 2.1. This is even more important when faced with a complex environment, where APIs and tools keep changing while the evaluation should maintain its consistency across time.
@@
Existing large-scale benchmarks may struggle to provide stable evaluations for various reasons. We propose several hypotheses for this issue. Firstly, the complexity of tasks involving tool usage makes it challenging for the common automatic evaluator, gpt-3.5, to function effectively as a discriminator. As discussed in Section 2.2, the evaluator cannot reliably determine whether a task is solvable or unsolvable, leading to variability in model performance due to this capability limitation. Secondly, the stability of API status for a significant portion of online tools (55.6% in ToolBench) is inconsistent. Users may be required to authorise the use of these tools or APIs, and tools provided by developers may be accessible during the initial construction of the benchmark but become unavailable later. This fluctuation further undermines the reliability and reproducibility of model performance assessments over time. This situation results in a problem where the constructed queries in the benchmarks may no longer be completed with their originally referenced tools. Consequently, it is crucial to strike a balance between enhancing the stability of these benchmarks and maintaining their diversity and scope.
@@
To address these issues, we propose a new benchmark named StableToolBench, which incorporates a virtual API system and a stable evaluation system. We first build a virtual API system to replace the real one. As a start, we build a caching system to store the outputs of API calls. This approach ensures the stability and reproducibility of API behaviours. Given the limited number of benchmark questions, our caching system can cover a significant number of API call scenarios. However, relying solely on a cache is insufficient because many APIs remain unavailable. To resolve this problem, we use large language models (LLMs) to simulate the behaviours of these APIs. Specifically, we feed the documentation and few-shot real API calls if available in the cache to LLMs and ask LLMs to mock the behaviour of the APIs given a request. As a result, users can always get responses from APIs in an indistinguishable way as long as the LLMs are accessible. On the whole, our system first tries to find a hit in the cache. Unless there is a cache miss and a real API call is not received, the simulated server will be used.
We then improve the evaluation system to make it more stable. We design two metrics (i.e., SoPR and SoWR) after judging solvable tasks and replace all the automatic evaluators with GPT-4 to mitigate the randomness and indistinguishability during evaluation. Experiments demonstrate that our virtual API system, when combined with the improved evaluation system, can provide stable evaluation against API modifications. Furthermore, our system exhibits significant reliability in terms of realism, diversity, and documentation following accuracy.
@@
The main contributions of our work are summarised as follows:
@@
  • Extensive experiments show that our benchmark provides much more stable model performance, robust to various types of API failures.
  • Besides enhanced stability, our virtual API system exhibits reality, diversity and reliability comparable to that of the real API system.
@@ 3.2 Stable Evaluation System
3.5 0613 (C) 52.2±$\pm{1.1}$ 47.3±$\pm{0.6}$ 53.6±$\pm{1.3}$ 42.5±$\pm{2.1}$ 35.8±$\pm{2.0}$ 48.1±$\pm{0.8}$ 46.6±$\pm{1.3}$
3.5 0613 (C) 60.3±$\pm{1.3}$ 66.2±$\pm{1.2}$ 67.1±$\pm{0.0}$ 59.1±$\pm{0.4}$ 51.3±$\pm{1.2}$ 73.8±$\pm{2.3}$ 63.0±$\pm{1.1}$
4 0613 (C) 45.5±$\pm{0.4}$ 57.4±$\pm{0.3}$ 48.8±$\pm{0.7}$ 43.0±$\pm{0.7}$ 46.5±$\pm{0.9}$ 48.1±$\pm{1.5}$ 48.2±$\pm{0.8}$
4 0613 (D) 57.3±$\pm{0.6}$ 57.3±$\pm{0.3}$ 60.9±$\pm{1.0}$ 57.9±$\pm{1.0}$ 51.3±$\pm{0.8}$ 66.4±$\pm{2.4}$ 58.5±$\pm{1.0}$
T-LLaMA (C) 32.3±$\pm{1.0}$ 40.3±$\pm{0.8}$ 36.7±$\pm{0.5}$ 34.7±$\pm{0.7}$ 25.2±$\pm{0.4}$ 33.9±$\pm{1.5}$ 33.9±$\pm{0.8}$
T-LLaMA (D) 44.5±$\pm{0.9}$ 49.6±$\pm{1.3}$ 48.9±$\pm{2.7}$ 50.8±$\pm{1.1}$ 31.9±$\pm{1.9}$ 53.6±$\pm{2.0}$ 46.6±$\pm{1.7}$
3.5 1106 (C) $\pm{0.5}$ 45.1±$\pm{1.4}$ 50.8±$\pm{0.3}$ 48.7±$\pm{0.8}$ 42.1±$\pm{0.4}$ 55.7±$\pm{0.0}$ 48.8±$\pm{0.6}$
3.5 1106 (D) 62.8±$\pm{0.3}$ 63.9±$\pm{1.2}$ 65.6±$\pm{0.3}$ 56.5±$\pm{0.7}$ 56.9±$\pm{1.2}$ 67.2±$\pm{1.3}$ 62.2±$\pm{0.8}$
4 Turbo (C) 52.8±$\pm{1.3}$ 56.6±$\pm{0.9}$ 51.9±$\pm{0.5}$ 51.9±$\pm{1.0}$ 52.8±$\pm{0.8}$ 52.5±$\pm{0.0}$ 53.1±$\pm{0.8}$
4 Turbo (D) 59.2±$\pm{0.5}$ 61.7±$\pm{0.7}$ 65.7±$\pm{1.0}$ 55.6±$\pm{0.6}$ 55.2±$\pm{0.4}$ 66.1±$\pm{4.3}$ 60.6±$\pm{1.3}$
Table 4: Solvable pass rate scores. We run all models once, evaluate three times and take the average results. “3.5 0613”, “4 0613”, “3.5 1106”, “4 Turbo”, “T-LLaMA” stands for gpt-3.5-turbo-0613, gpt-4-0613, gpt-3.5-turbo-1106, gpt-4-turbo-preview, ToolLLaMA v2 respectively. C and D stand for CoT and DFS respectively. The experiments below follow the denotation. We use gpt-4-turbo-2024-04-09 as the evaluator. Evaluation done on May 2024.
@@
We then report Solvable Pass Rate (SoPR) and Solvable Win Rate (SoWR) based on our obtained solvable tasks. Due to the limitation of gpt-3.5-turbo-16k in tool learning, we uniformly adopt gpt-4-turbo as the automatic evaluator. SoPR is in essence PR with all tasks solvable and only assesses the answers using the same prompt in ToolBench. The evaluator assigns outcomes of answers categorised as Solved, Unsolved, or Unsure, which respectively contribute scores of 1, 0.5, and 0 to the overall SoPR calculation. As for SoWR, when one is solved and the other is unsolved, the solved one wins. Under other circumstances, gpt-4-turbo will be used to make a win-lose decision.
@@ 4.1 Performance
3.5 0613 (D) 60.7 67.3 59.5 63.2 62.1 75.4 64.7
4 0613 (C) 54.6 58.8 58.2 75.5 60.5 62.3 61.7
4 0613 (D) 62.6 62.7 58.2 74.5 62.9 67.2 64.7
T-LLaMA (C) 31.3 28.1 33.5 35.8 33.9 24.6 31.2
T-LLaMA (D) 44.8 45.8 44.3 59.4 41.1 50.8 47.7
3.5 1106 (C) 47.2 47.7 44.9 50.9 54.0 62.3 51.2
3.5 1106 (D) 55.8 53.6 51.9 68.9 59.7 68.9 59.8
4 Turbo (C) 71.2 77.1 61.4 79.2 71.8 67.2 71.3
4 Turbo (D) 73.0 75.2 68.4 77.4 66.9 60.7 70.2
Table 5: Solvable Win Rate scores. We run all models once against GPT-3.5-Turbo-0613 + CoT and evaluate three times. We follow the ToolBench implementation to take the most frequent result for each query during evaluation. The experiments below follow the denotation. We use gpt-4-turbo-2024-04-09 as the evaluator. Evaluation done on May 2024.
@@ 4.2 Stability of Virtual API Server
Figure 7: Performance change when manually making APIs down with our virtual online API system. The results are averaged over all six groups. Solving rates are reported. We run each experiment one time and evaluate it three times and take the average score. Unless otherwise stated, gpt-4-turbo-preview at the time of testing is used in this section. This experiment was done in Feb 2024.
@@ 4.3 Turing Test of API Simulator
In addition to the Turing Test mentioned above, we assess the quality of LLM simulations by evaluating the adherence of simulated outputs to their corresponding documentation. To conduct this evaluation, we randomly sample 50 simulated outputs along with their documentation from the Turing Test dataset. A human evaluator is then tasked with determining whether the LLM simulations reasonably follow the provided documentation. The results indicate that 90% of the simulations are deemed reasonable, 6% are considered unreasonable, and 4% are uncertain. These findings suggest that the LLM is highly capable of generating simulated responses that adhere closely to the provided documentation.
@@ 4.5 Effectiveness of Caching System
4.6 Human Evaluation of Evaluator
Considering that GPT-3.5 is limited to evaluating the performance in tool learning, we replace the automatic evaluators with stronger LLMs. In this section, we manually assess the correctness of different automatic evaluators. Here, we sample 100 task-solvable questions, 50 answer-solving questions in the PR / SoPR evaluation, and 50 comparison questions in the WR / SoWR evaluation from the experiments running.
@@ 4.6 Human Evaluation of Evaluator
Table 7: Cache hit rate (%) with various models and methods. Final, Mid, and Start represent the final version of the cache, the mid-way version containing 151,152 (91.6%) items of the final version, and the starting version containing only the train and test set. Experiments are independent runs of Section 4.1 with fixed cache, run on 13 Feb 2024.
We then collect all the corresponding answers of different LLMs during previous evaluations. These questions are further manually labeled by three human annotators to obtain the ground truth. With the ground truth, we calculate the accuracy scores of these models and show the results in Table 8. It can be seen that our used LLMs (i.e., Claude 2, Gemini, and GPT-4) are much better than GPT-3.5 in ToolBench to determine the solvability of tasks, where the Gemini and GPT-4 outperform by a large margin. In both the evaluation of answers and comparison, GPT-4 significantly outperforms GPT-3.5, especially in the comparison to compute WR. It is worth noting that all the accuracies of GPT-3.5 are lower than 70%, indicating that GPT-3.5 cannot assess the performance in tool learning.
@@ 5 Related Work
Recent studies have shed light on the burgeoning capabilities of LLMs in understanding and mastering tools (Li et al., 2023a; Patil et al., 2023; Yang et al., 2023b; Song et al., 2023; Tang et al., 2023; Ye et al., 2024; Xu et al., 2023). Gaining access to external tools endows LLMs with real-time factual knowledge (Yang et al., 2023a), multimodal functionalities (Gupta and Kembhavi, 2023), and specialised skills in vertical domains (Jin et al., 2024). However, few work has been done to explore the stability of the tool environment in specific benchmarks and how it affects the LLMs’ performance in tool-augmented tasks.
@@
Evaluating the performance of LLMs in tool-augmented tasks presents unique challenges and opportunities. Numerous works have been developed for the assessment of tool utilisation, primarily emphasising response comparison  (Zhuang et al., 2023), tool call accuracy  (Patil et al., 2023), or a synthesis of these aspects  (Li et al., 2023a). Distinguishing itself, Qin et al. (2023c) introduces an innovative methodology by integrating a large language model (LLM) as a judge to evaluate the comprehensive solution path. Subsequent research  (Wang et al., 2023) focuses on the multi-turn interaction capabilities of LLMs with both tools and user feedback. In a departure from the aforementioned approaches, Chen et al. (2023b) presents itself as the inaugural benchmark specifically tailored for the fine-grained assessment of tool utilisation capabilities. However, there exists a gap in the literature concerning the exploration of evaluation stability when evaluating the tool usage capabilities of LLMs.